glizxzy:

On the surface, it seems easy. Can you think of the integers for x, y, and z so that x³+y³+z³=8? Sure. One answer is x = 1, y = -1, and z = 2. But what about the integers for x, y, and z so that x³+y³+z³=42?

glizxzy:

nobody knows the answer?

jhonyy9:

1^3 = ? and (-1)^3 = ?

jhonyy9:

and so in this way z need being what number ? to get 8 ?

glizxzy:

close 10^3 + 9^3 = 12^3

jhonyy9:

no

jhonyy9:

2^3 = 2*2*2

glizxzy:

dang your right

glizxzy:

how you do that

jhonyy9:

i ve to you above

jhonyy9:

i ve wrote

glizxzy:

ok what is our iq